40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5597     Entry time: Sun Oct 2 18:33:36 2011
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: inversion of the input matrix on ETMX, ITMX and PRM 

The input matrices of ETMX, ITMX and PRM have been newly inverted.

Those were the ones having some troubles (see #5444, #5547).

After a coarse adjustment of the damping Q factors, they look damping happily.

 

    optic                    spectra    inverted matrix   BADNESS
ETMX ETMX.png       pit     yaw     pos     side    butt
UL    0.848   0.108   1.578   0.431   1.025 
UR    0.182  -1.892   1.841  -0.128  -1.172 
LR   -1.818  -1.948   0.422  -0.122   0.939 
LL   -1.152   0.052   0.159   0.438  -0.864 
SD    1.756  -3.794  -0.787   1.000  -0.132 
  5.37028
ITMX ITMX.png     pit     yaw     pos     side    butt
UL    0.510   0.584   1.228   0.458   0.203 
UR    0.783  -1.350   0.348  -0.050   0.555 
LR   -1.217   0.065   0.772   0.264   0.312 
LL   -1.490   2.000   1.652   0.772  -2.930 
SD   -0.635   0.853  -1.799   1.000  -1.597  
 7.5125
PRM PRM.png       pit     yaw     pos     side    butt
UL    0.695   1.422   1.774  -0.333   0.954 
UR    1.291  -0.578   0.674  -0.068  -0.939 
LR   -0.709  -1.022   0.226   0.014   0.867 
LL   -1.305   0.978   1.326  -0.251  -1.240 
SD    0.392  -0.437  -0.500   1.000   0.420
 5.09674

 

(some notes)

 Before running the peakFit scripts, I woke up the nds2 sever process on Mafalda because it hasn't been recovered from the power outage.

To start the nds2 process I followed the instruction by Jamie (#5094).

Then I started requesting the data of the last night's free swinging test (#5594)

However the NDS2_GetData command failed everytime when data with long duration were requested.

It maybe because some of the data are missing in sometimes, but I haven't seriously checked the data stored in fb.

So for the reason, I had to use a short duration of 1200 sec (default is 3600 sec). That's why spectra look noisier than usual.

ELOG V3.1.3-